The largest number which always divides the sum of any pair of consecutive odd numbers is 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

1 + 3 = 4 = 4/4 = 1
3 + 5 = 8 = 8/4 = 2
The largest number which always divides the sum of any pair of consecutive oddnumbers is 4.

Thanks